OpenStudy (anonymous):

find the exact area of a circle having the given circumference: 3pi

OpenStudy (anonymous):

\[3\pi\]

OpenStudy (campbell_st):

well start with the circumference the circumference of a circle in terms of radius is \[C = 2 \pi r\] and you know \[C = 3\pi \] so \[3\pi = 2\pi r\] you need to solve for r...can you do that

OpenStudy (anonymous):

so 3* 3.14? =9.42 then what?

OpenStudy (campbell_st):

it asks for an exact value so just use p to find r if \[3\pi = 2 \pi r\] then \[r = \frac{3\pi}{2\pi}\] eliminate the common factor to get \[r = \frac{3}{2}\] the exact value means you need to use fractions and pi... no decimals... can you now find the area of the circle.?

OpenStudy (anonymous):

\[\frac{ 3 }{ 2 } X \frac{ 3 }{ 2 } X 3.14 =A ?\]

OpenStudy (anonymous):

9/4 pi

OpenStudy (campbell_st):

correct